OpenStudy (anonymous):

HELP ! TOTALLY CONFUSED !! Simplify –(3ab^2)^–3

OpenStudy (anonymous):

\[x^{(-n)} = \frac{ 1 }{ x^{n} }\]

OpenStudy (anonymous):

what ? o.O

OpenStudy (anonymous):

this is \[\frac{ -1 }{(3ab^{2})^{3} }\] it can be simplified more though

OpenStudy (anonymous):

use law of exponents

OpenStudy (anonymous):

would it be : \[\frac{ -1 }{ 27a^3b^6 }\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thanks ! :D